# Learning Enhanced Ensemble Filters

### v1

@@ 4.2.3 Learning the Localization
Localization is implemented by damping covariances in the empirical covariance matrices based on distance. This is typically achieved by applying a Hadamard product (elementwise product, denoted by $\circ$) between the empirical covariance matrix and a predefined localization weight matrix $L$. For the dynamic model (2.1) with the state in $d_{v}$ dimension space (i.e. $v\in\mathbb{R}^{d_{v}}$), $L\in\mathbb{R}^{d_{v}\times d_{v}}$ is calculated by
@@
where $g:\mathbb{R}\rightarrow\mathbb{R}$ is a function that maps distance values to localization weights, and $D_{k\ell}$ is the distance between index $k$ and $\ell$. To incorporate the localization technique into EnKF, the Kalman gain formula (2.11b) is modified as
@@
$K_{\theta}=K_{\theta}^{(1)}\circ L_{\theta}^{(1)}\left(K_{\theta}^{(2)}\circ L_{\theta}^{(2)}+\Gamma\right)^{-1}.$ (4.35)
The localization weight matrices $L_{\theta}^{(1)}$ and $L_{\theta}^{(2)}$ are learned, and we now detail how this is achieved. As in the EnkF, where localization is defined using (4.33), the core idea is to learn a parameterized distance-to-weight function $g_{\theta}:\mathbb{R}\rightarrow\mathbb{R}$.
In practice, it is not necessary to learn the entire function $g_{\theta}$ since we only have finite number of different distances, defined by the vector of grid values, for discretized PDEs, and directly for vector indices in the setting of ODEs. We assume that the unique distance values in $\{D_{k\ell}\}_{k,\ell=1}^{d_{v}}$ are denoted as $\{D_{1},D_{2},\ldots,D_{N_{D}}\}$, where $N_{D}$ is the number of distinct distances. Given a vector $\widehat{g}=[\widehat{g}_{1},\widehat{g}_{2},\ldots,\widehat{g}_{N_{D}}]\in\mathbb{R}^{N_{D}}$, we can define function $g_{\theta}$ by
@@
The parameterized localization weight matrices $L_{\theta}^{(1)}$ and $L_{\theta}^{(2)}$ are calculated based on this $g_{\theta}$. In B.1, we provide a step-by-step example illustrating the computation of the parameterized localization weight matrices.
@@
  • The final layer of the $F^{\mathrm{loc}}$ is a $2\times\mathrm{softmax}$ function, ensuring all outputs lie within the interval $[0,2]$. Although amplifying underestimated covariances is typically the responsibility of inflation, our architecture jointly learns both localization and inflation. Thus, we allow the weights to range in $[0,2]$ rather than $[0,1]$ to grant additional flexibility to the localization component.
  • Instead of learning a continuous parametric form for $g_{\theta}$, we learn its values only at discrete distances $\{D_{i}\}_{i=1}^{N_{D}}$. This reduces model complexity while preserving localization flexibility.
  • While EnKF-based localization methods typically use fixed weight matrices across time steps (e.g., using the Gaspari–Cohn function [45]), there is much interest in adaptive approaches [46]. Our learning-based approach adapts these matrices for each time step $j$ using neural networks. As shown in (4.37) (time index $j$ omitted for simplicity), the network processes time-step-specific inputs, allowing the localization scheme to dynamically adjust based on the current system state.
@@ 4.3 Loss Function
Our training is conducted for fixed dynamics $\Psi$, observation operator $h$, and noise covariance matrices $\Sigma$ and $\Gamma$. The time step $\Delta t$ is also fixed (and incorporated in $\Psi$). Given choice of hyperparameters $M,J\in\mathbb{N}$, the training data consists of $M$ trajectories of length $J+1$, extracted from a single long trajectory generated by propagating the dynamic model (2.1) for $M(J+1)$ steps. Each sub-trajectory $\{v^{\dagger}_{j}\}_{j=0}^{J}$ has corresponding observations $\{y^{\dagger}_{j}\}_{j=1}^{J}$ generated according to (2.2).
For each trajectory, we generate ensemble-based data assimilation outputs to compare with the observations from the state observation model, enabling us to train the parameters $\theta_{\mathrm{ST}}$, $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$, for the set transformer $F^{\mathrm{ST}}$ (4.25), and MLPs $F^{\mathrm{gain}}$ (4.28), $F^{\mathrm{infl}}$ (4.31) and $F^{\mathrm{loc}}$ (4.37) respectively. We initialize an ensemble $\{v_{0}^{(n)}\}_{n=1}^{N}$ at the initial time step $j=0$ for each sub-trajectory, with members sampled i.i.d. from a Gaussian distribution:
@@
where $(K_{\theta})_{j+1}$ is our learned parameterized gain matrix, including localization, and $\widehat{u}^{(n)}_{\theta}$ is the inflation term according to Subsection 4.2. For all $\theta$ we choose $v_{0}^{(n)}(\theta)=v_{0}^{(n)}$. The ensembles $\{v_{j}^{(n)}(\theta)\}_{n=1}^{N}$ depend on the trainable parameters $\theta$ for $j=1,2,\ldots,J$.
@@
$\bar{v}_{j}(\theta)=\frac{1}{N}\sum_{n=1}^{N}v_{j}^{(n)}(\theta).$ (4.40)
To train our neural network parameters effectively, we introduce a loss function that quantifies the discrepancy between the estimated states and the true states across a given sub-trajectory $V=\{v^{\dagger}_{j}\}_{j=0}^{J}$, given by:
$\mathcal{L}(\theta)=\frac{1}{J}\sum_{j=1}^{J}\frac{\|\bar{v}_{j}(\theta)-v_{j}^{\dagger}\|_{2}^{2}}{\|v_{j}^{\dagger}\|_{2}^{2}},$ (4.41)
which measures the relative accuracy of our ensemble mean predictions compared to the true trajectory. Our loss (4.41) is a relative squared L2 loss, which is preferred over standard L2 loss for two main reasons. First, using squared L2 norm avoids computing square roots, which can lead to numerical instabilities when calculating gradients, especially when the error approaches zero. Second, normalizing by the true state magnitude ($\|v_{j}^{\dagger}\|_{2}^{2}$) makes the loss scale-invariant across different state variables and trajectory segments, ensuring balanced parameter updates regardless of the absolute magnitudes of the state components. This relative formulation is particularly important in dynamical systems where state variables may span different orders of magnitude.
Since we have $M$ training sub-trajectories we denote the loss corresponding to the $m^{th}$ trajectory by $\mathcal{L}_{m}$. In the training process, when we choose a minibatch with the index set $M_{B}\subset[M]$, the minibatch loss is
$\mathcal{L}_{M_{B}}(\theta)=\frac{1}{|M_{B}|}\sum_{m\in M_{B}}\mathcal{L}_{m}(\theta).$ (4.42)
In our training framework, the parameter $\theta$ is optimized based on the above-defined loss $\mathcal{L}_{M_{B}}$.
@@
Computing the loss (4.41) requires evaluating ensemble estimations $\{\{v_{j}^{(n)}\}_{n=1}^{N}\}_{j=1}^{J}$ across the entire trajectory. For the ensemble at step $j$, this involves nested calls to $K_{\theta}$ up to $j$ times. These nested calls arise from the sequential nature of the trajectory, where each step depends on the evaluation of the Kalman gain from the previous steps, requiring backpropagation through the entire computational history. With large trajectory lengths $J$, this causes slow training and numerical instability since it is nearly equivalent to training a very deep neural network.
To address this issue, we introduce a gradient-detach hyperparameter $J_{0}\in\mathbb{N}$. This hyperparameter limits the gradient computation for the loss related to time step $j$ to only the steps $j,j-1,\ldots,j-J_{0}+1$ This approach assumes that the trajectory at time step $j$ is not significantly affected by parameter changes in the Kalman gain for earlier time steps, which is reasonable given the diminishing influence of distant past states in many dynamical systems. This technique significantly improves training speed and stability.
@@
While this work focuses on state estimation using relative squared L2 loss (4.41), a promising direction for future work is to consider probability-based loss functions that address the complete filtering problem. Such approaches would facilitate the learning of data assimialtion algorithms tuned to accurately approximate the filtering distribution $\mathbb{P}(v_{j}|Y_{j})$; in contrast the current work simply tunes to the point estimator defined by the mean.
@@ 4.4 Fine-Tuning Inflation & Localization
Following the developments in Subsection 4.2, we may apply a set transformer (4.25) to process ensembles of any size into fixed-dimensional feature vectors. This allows for pretraining our MNMEF with ensemble size $N$ and inference with $N^{\prime}\neq N$. However in classical EnKF formulations, hyperparameters such as inflation and localization depend on the ensemble size, and allowing for this dependence significantly enhances performance. Thus we consider fine-tuning a subset of the parameters for inference at ensemble size $N^{\prime}\neq N$. We now detail this methodology.
The architecture defined in Subsection 4.2 is based on trainable parameters $\theta_{\mathrm{ST}}$, for the set transformer $F^{\mathrm{ST}}$ (4.25), and $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$ for the MLPs $F^{\mathrm{gain}}$ (4.28), $F^{\mathrm{infl}}$ (4.31) and $F^{\mathrm{loc}}$ (4.37) respectively. For efficient fine-tuning when transitioning from ensemble size $N$ to $N^{\prime}$, we freeze $\theta_{\mathrm{ST}}$ and only fine-tune $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$. This specific choice of fine-tuning is effective because: (1) freezing $\theta_{\mathrm{ST}}$ significantly reduces computational costs, as the set transformer contains the majority of network parameters; (2) the remaining parameters, though small in number, measurably improve performance if adapated to the ensemble size.
Recall the training loss defined by (4.40), (4.41),
$\mathcal{L}^{(N)}(\theta_{\mathrm{ST}},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}})=\frac{1}{J}\sum_{j=1}^{J}\frac{\|\frac{1}{N}\sum_{n=1}^{N}v_{j}^{(n)}(\theta)-v_{j}^{\dagger}\|_{2}^{2}}{\|v_{j}^{\dagger}\|_{2}^{2}},$ (4.43)
where the superscript $N$ emphasizes dependence on the ensemble size. In pretraining with size $N$ we solve
$\theta_{\mathrm{ST}}^{(N)},\theta_{\mathrm{gain}}^{(N)},\theta_{\mathrm{infl}}^{(N)},\theta_{\mathrm{loc}}^{(N)}=\operatorname*{arg\,min}\;\mathcal{L}^{(N)}(\theta_{\mathrm{ST}},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}}).$ (4.44)
@@
$\theta_{\mathrm{gain}}^{(N^{\prime})},\theta_{\mathrm{infl}}^{(N^{\prime})},\theta_{\mathrm{loc}}^{(N^{\prime})}=\operatorname*{arg\,min}_{\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}}}\;\mathcal{L}^{(N^{\prime})}(\theta_{\mathrm{ST}}^{(N)},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}})$ (4.45)
with fixed $\theta_{\mathrm{ST}}^{(N)}$. For multiple trajectories, the loss is averaged across them.
@@
Although $\theta_{\mathrm{ST}}^{(N)}$ remains unchanged after the fine-tuning, the feature vector $f_{v}^{(N^{\prime})}$ differs from pretraining $f_{v}$ since the input ensemble size changes from $N$ to $N^{\prime}$.
@@ 5 Numerical Experiments
In this section we present numerical experiments to demonstrate the superiority of our proposed method, MNMEF, in comparison with state of the art existing ensemble methods; the experiments also serve to validate the effectiveness of several specific details within our approach. First, in Subsections 5.2, 5.3, and 5.4, we compare our method with several classical approaches across different datasets. In Subsection 5.6, we conduct experiments to study the learning-based inflation and localization methods proposed in Subsection 4.2.3. Subsequently, in Subsection 5.7, we perform experiments to illustrate the efficiency and effectiveness of our proposed fine-tuning method, introduced in Subsection 4.4. The code for our method is publicly available  [^4] https://github.com/wispcarey/DALearning and can be used to reproduce all our experiments.

### v2

@@ 4.2.3 Learning the Localization
Localization is implemented by damping covariances in an empirical covariance matrix based on distance. This is typically achieved by applying a Hadamard product (elementwise product, denoted by $\circ$) between the empirical covariance matrix and a predefined localization weight matrix $L$. For the dynamic model (2.1) with the state in $d_{v}$-dimensional space (i.e., $v\in\mathbb{R}^{d_{v}}$), $L\in\mathbb{R}^{d_{v}\times d_{v}}$ is calculated by
@@
where $g:\mathbb{R}\rightarrow\mathbb{R}$ is a function that maps distance values to localization weights (e.g., the Gaspari–Cohn function [45]), and $D_{k\ell}$ is the distance between index $k$ and $\ell$. To incorporate the localization technique in the EnKF, the Kalman gain formula (2.11b) is modified as
@@
$K_{\theta}=K_{\theta}^{(1)}\circ L_{\theta}^{(1)}\left(K_{\theta}^{(2)}\circ L_{\theta}^{(2)}+\Gamma\right)^{-1},$ (4.34)
The localization weight matrices $L_{\theta}^{(1)}$ and $L_{\theta}^{(2)}$ follow the structure in (4.32), where $g$ is replaced by a parameterized distance-to-weight function $g_{\theta}:\mathbb{R}\rightarrow\mathbb{R}$.
In practice, it is not necessary to learn the entire function $g_{\theta}$ since we only have a finite number of different distances, defined by the spatial grid for discretized PDEs, and time indices in the setting of ODEs. We denote the unique distance values in $\{D_{k\ell}\}_{k,\ell=1}^{d_{v}}$ as $\{D_{1},D_{2},\ldots,D_{N_{D}}\}$, where $N_{D}$ is the number of distinct distances. Given a vector $\widehat{g}=[\widehat{g}_{1},\widehat{g}_{2},\ldots,\widehat{g}_{N_{D}}]\in\mathbb{R}^{N_{D}}$, we can define the function $g_{\theta}$ by
@@
The parameterized localization weight matrices $L_{\theta}^{(1)}$ and $L_{\theta}^{(2)}$ are calculated based on the output of this learned function $g_{\theta}$. In C.2, we provide a step-by-step example illustrating the computation of the parameterized localization weight matrices.
@@
  • The final layer of $F^{\mathrm{loc}}$ is a $\mathrm{softmax}$ function scaled by 2, ensuring all outputs lie within the interval $[0,2]$. Although amplifying underestimated covariances is typically the responsibility of inflation alone, our architecture jointly performs localization and inflation. Thus, we allow the weights to range in $[0,2]$ rather than $[0,1]$ to grant additional flexibility in the localization component.
  • Instead of learning a continuous parametric form for $g_{\theta}$, we learn its values only at discrete distances $\{D_{i}\}_{i=1}^{N_{D}}$. This reduces model complexity while preserving localization expressivity.
  • While EnKF-based localization methods typically use fixed weight matrices across time steps (e.g., using the Gaspari–Cohn function [45]), there is much interest in adaptive approaches [46]. Our learning-based approach adapts these matrices for each time step $j$ using neural networks via its dependence on the ensemble. As shown in (4.36) (time index $j$ omitted for simplicity), the network processes time-step-specific inputs, allowing the localization scheme to dynamically adjust based on the current system state.
@@ 4.3 Loss Function
Our training is conducted for fixed dynamics $\Psi$, observation operator $h$, and noise covariance matrices $\Sigma$ and $\Gamma$. The time step $\Delta t$ is also fixed (and incorporated in $\Psi$). The initial condition is a Gaussian with a fixed covariance $C_{0}$, i.e. $v_{0}\sim\mathsf{N}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},C_{0})$. For some $M,J\in\mathbb{N}$, the training data consists of $M$ trajectories of length $J+1$, extracted from a single long trajectory generated by propagating the dynamic model (2.1) for $M(J+1)$ steps. Each sub-trajectory $\{v^{\dagger}_{j}\}_{j=0}^{J}$ has corresponding observations $\{y^{\dagger}_{j}\}_{j=1}^{J}$ generated according to (2.2).
For each trajectory, we generate estimated states from our ensemble-based data assimilation model to compare with the true states from the dynamic model, enabling us to train the parameters $\theta_{\mathrm{ST}}$, $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$, for the set transformer $F^{\mathrm{ST}}$ (4.25), and MLPs $F^{\mathrm{gain}}$ (4.28), $F^{\mathrm{infl}}$ (4.30) and $F^{\mathrm{loc}}$ (4.36) respectively.
We initialize an ensemble $\{v_{0}^{(n)}\}_{n=1}^{N}$ at the initial time step $j=0$ for each sub-trajectory, with members sampled i.i.d. from a Gaussian distribution with the mean $v^{(n)}_{0}$ and the covariance $C_{0}$:
@@
where $(K_{\theta})_{j+1}$ is our learned parameterized gain matrix, including localization, and $\widehat{u}^{(n)}_{\theta}$ is the inflation term according to Subsection 4.2. For all parameters $\theta$ we choose $v_{0}^{(n)}(\theta)=v_{0}^{(n)}$. The ensembles $\{v_{j}^{(n)}(\theta)\}_{n=1}^{N}$ depend on the trainable parameters $\theta$ for $j=1,2,\ldots,J$.
@@
$\overline{v}_{j}(\theta)=\frac{1}{N}\sum_{n=1}^{N}v_{j}^{(n)}(\theta).$ (4.39)
To train our neural network parameters effectively, we introduce a loss function that quantifies the discrepancy between the estimated states and the true states across the $m$-th sub-trajectory $V_{m}=\{v^{\dagger}_{j}\}_{j=1}^{J}$, given by:
$\mathcal{L}_{m}(\theta)=\frac{1}{J}\sum_{j=1}^{J}\frac{\|\overline{v}_{j}(\theta)-v_{j}^{\dagger}\|_{2}^{2}}{\|v_{j}^{\dagger}\|_{2}^{2}},$ (4.40)
which measures the relative accuracy of our ensemble mean predictions compared to the true trajectory. Our loss (4.40) is a relative squared $L_{2}$ loss, which is preferred over the standard $L_{2}$ loss for two main reasons. First, using the squared $L_{2}$ norm avoids computing square roots, which can lead to numerical instabilities when calculating gradients, especially when the error approaches zero. Second, normalizing by the true state magnitude ($\|v_{j}^{\dagger}\|_{2}^{2}$) makes the loss scale-invariant across different state variables and trajectory segments, ensuring balanced parameter updates regardless of the absolute magnitudes of the state components. This relative formulation is particularly important in dynamical systems where state variables may span different orders of magnitude.
In the training process, the loss is averaged across $M$ training trajectories indexed by $m\in[M]$,
$\mathcal{L}_{[M]}(\theta)=\frac{1}{M}\sum_{m\in[M]}\mathcal{L}_{m}(\theta).$ (4.41)
The trainable parameter $\theta$ is optimized based on the above-defined loss $\mathcal{L}_{[M]}$. In practice, the loss $\mathcal{L}_{[M]}$ is approximated by a mini-batch as a subset of $[M]$.
@@
Computing the loss (4.40) requires evaluating ensemble estimates $\{\{v_{j}^{(n)}(\theta)\}_{n=1}^{N}\}_{j=1}^{J}$ across the entire trajectory. For the ensemble at step $j$, this involves nested evaluations of $K_{\theta}$ up to $j$ times. These nested evaluation arise from the sequential nature of the trajectory, where each step depends on the computing of the Kalman gain from the previous steps, requiring backpropagation through the entire computational history. With large trajectory lengths $J$, this causes slow training and numerical instability since it is nearly equivalent to training a very deep neural network.
To address this issue, we introduce a gradient-detach hyperparameter $J_{0}\in\mathbb{N}$. This hyperparameter limits the gradient computation for the loss related to time step $j$ to only the steps $j,j-1,\ldots,j-J_{0}+1$. This approach assumes that the trajectory at time step $j$ is not significantly affected by parameter changes in the Kalman gain for earlier time steps, which is reasonable given the diminishing influence of distant past states in many dynamical systems. This technique significantly improves training speed and stability.
@@
While this work focuses on state estimation using the relative squared $L_{2}$ loss (4.40), a promising direction for future work is to consider probabilistic-based loss functions that address the complete filtering problem. Such approaches would facilitate the learning of data assimilation algorithms tuned to accurately approximate the filtering distribution $\mathbb{P}(v_{j}|Y_{j})$; in contrast the current work simply tunes to the point estimator defined by the mean.
@@ 4.4 Fine-Tuning Inflation & Localization
Following the developments in Subsection 4.2, we may apply a set transformer (4.25) to process ensembles of any size into fixed-dimensional feature vectors. This allows for pretraining our MNMEF with ensemble size $N$ and perform inference with $N^{\prime}\neq N$. However in classical EnKF formulations, hyperparameters such as inflation and localization depend on the ensemble size, and allowing for this dependence significantly enhances performance. Thus, we consider fine-tuning a subset of the parameters for inference at ensemble size $N^{\prime}\neq N$. We now detail this methodology.
The architecture defined in Subsection 4.2 is based on trainable parameters $\theta_{\mathrm{ST}}$, for the set transformer $F^{\mathrm{ST}}$ (4.25), and $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$ for the MLPs $F^{\mathrm{gain}}$ (4.28), $F^{\mathrm{infl}}$ (4.30) and $F^{\mathrm{loc}}$ (4.36) respectively. For efficient fine-tuning when transitioning from ensemble size $N$ to $N^{\prime}$, we freeze $\theta_{\mathrm{ST}}$ and only fine-tune $\theta_{\mathrm{gain}}$, $\theta_{\mathrm{infl}}$, and $\theta_{\mathrm{loc}}$. This specific choice of fine-tuning is effective because: (1) freezing $\theta_{\mathrm{ST}}$ significantly reduces computational costs, as the set transformer contains the majority of network parameters; (2) the remaining parameters, though small in number, measurably improve performance when they are adapted to the ensemble size.
The training loss for the $m$-th sub-trajectory $V_{m}=\{v^{\dagger}_{j}\}_{j=1}^{J}$ is defined by (4.39), (4.40):
$\mathcal{L}_{m}^{(N)}(\theta_{\mathrm{ST}},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}})=\frac{1}{J}\sum_{j=1}^{J}\frac{\|\overline{v}_{j}(\theta)-v_{j}^{\dagger}\|_{2}^{2}}{\|v_{j}^{\dagger}\|_{2}^{2}},$ (4.42)
where we add the superscript $N$ to emphasize dependence on the ensemble size. In pretraining with size $N$ we solve
$\theta_{\mathrm{ST}}^{(N)},\theta_{\mathrm{gain}}^{(N)},\theta_{\mathrm{infl}}^{(N)},\theta_{\mathrm{loc}}^{(N)}=\operatorname*{arg\,min}_{\theta_{\mathrm{ST}},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}}}\;\mathcal{L}_{[M]}^{(N)}(\theta_{\mathrm{ST}},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}}).$ (4.43)
@@
$\theta_{\mathrm{gain}}^{(N^{\prime})},\theta_{\mathrm{infl}}^{(N^{\prime})},\theta_{\mathrm{loc}}^{(N^{\prime})}=\operatorname*{arg\,min}_{\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}}}\;\mathcal{L}^{(N^{\prime})}_{[M]}(\theta_{\mathrm{ST}}^{(N)},\theta_{\mathrm{gain}},\theta_{\mathrm{infl}},\theta_{\mathrm{loc}})$ (4.44)
with fixed $\theta_{\mathrm{ST}}^{(N)}$. In practice, we use small mini-batches of $[M]$ to estimate the losses in (4.43) and (4.44).
@@
Although $\theta_{\mathrm{ST}}^{(N)}$ remains unchanged after the fine-tuning, the feature vector $f_{v}^{(N^{\prime})}$ differs from the pretraining features $f_{v}$ since the input ensemble size changes from $N$ to $N^{\prime}$.
@@ 5 Numerical Experiments
In this section we present numerical experiments to demonstrate the improved numerical benefit of our proposed method, MNMEF, in comparison with leading existing ensemble methods; the experiments also serve to validate the effectiveness of several specific details within our approach. First, in Subsections 5.2, 5.3, and 5.4, we compare our method with several classical approaches across different dynamical systems. In Subsection 5.5, we show the robustness of our method to the inherit randomness in test trajectories. In Subsection 5.6, we conduct experiments to study the learning-based inflation and localization methods proposed in Subsections 4.2.2 and 4.2.3. Subsequently, in Subsection 5.7, we perform experiments to illustrate the efficiency and effectiveness of our proposed fine-tuning method, introduced in Subsection 4.4. The code for our method is publicly available [^4] https://github.com/wispcarey/DALearning and can be used to reproduce all of our experiments.